If a product normally retails for $40.00 and we give you a $10 discount, what percentage discount do you receive?

Answer by stanbon(75887)

If a product normally retails for $40.00 and we give you a $10 discount, what percentage discount do you receive?
-------------
Amount of discount = $10
Percent of discount:10/40 = 1/4 = 25%
------------------------------------------
Cheers,
Stan H.